# Free multiderivations of connected subgraph arrangements

### v2

@@ 5. Classifying free multiplicities on connected subgraph arrangements
Since connected subgraph arrangements are locally $A_{2}$ and admit a positive system of defining equations, the extension techniques introduced by Yoshinaga in [Y09] are well suited to attack the problem of classifying free multiplicities on connected subgraph arrangements and their subarrangements.
We define a subarrangement of ${\mathscr{A}}_{C_{3}}$ for which we give a complete list of free multiplicities. This classification was achieved previously by DiPasquale and Wakefield in [DW18], where this arrangement appeared in a different context, using homological algebra methods.
@@
We define the deleted $C_{3}$ arrangement $\mathscr{D}$ as $\mathscr{D}:={\mathscr{A}}_{C_{3}}\backslash\{H_{123}\}$. The defining polynomial of $\mathscr{D}$ is given by
@@
    If $\mu=(k,1,k,2k-4)$ or $\mu=(k,2k-4,k,1)$, then $({\mathscr{A}},\mu)$ is free with exponents $(2k-2,2k-1)$.
@@
    If $\mu=(k+1,1,k+1,2k-4)$ or $\mu=(k+1,2k-4,k+1,1)$, then $({\mathscr{A}},\mu)$ is free with exponents $(2k-1,2k)$.
@@
(3) and (4): Start with $\mu^{\prime}=(k,0,k,2k-4)$ resp. $\mu^{\prime}=(k+1,0,k+1,2k-4)$ and use Theorem 2.16 to see that $({\mathscr{A}},\mu^{\prime})$ is free with exponents $(2k-2,2k-2)$ resp. $(2k-1,2k-1)$. So $({\mathscr{A}},\mu)$ has to be free with exponents $(2k-2,2k-1)$ resp. $(2k-1,2k)$.
@@ 5.2. Non-Free multiplicities on $\mathscr{D}$
For the forward implication assume that $(\mathscr{D},\mu)$ is free. For some $X\in L({\mathscr{A}})_{2}$ and the corresponding rank $2$ localization ${\mathscr{A}}_{X}=\{H_{1},H_{2},H_{3}\}$ or ${\mathscr{A}}_{X}=\{H_{1},H_{2}\}$ which is free with $\exp({\mathscr{A}}_{X},\mu_{X})=(d_{1}^{X},d_{2}^{X})$, we write its summand $d_{1}^{X}d_{2}^{X}$ in $\operatorname{LMP}(2)$ as $[\mu(H_{1}),\mu(H_{2}),\mu(H_{3})]$ or $[\mu(H_{1}),\mu(H_{2})]$ respectively. Using this notation, $\operatorname{LMP}(2)$ of $(\mathscr{D},\mu)$ equals
@@
$a(b+1)+a(c+1)+b(c+1)+a+b+c+3=3a+2b+c+ab+ac+bc+3.$
Define the multiplicity $\mu_{1}=(a,b,c,0,0,0)$ on $\mathscr{D}$, then $(\mathscr{D},\mu_{1})$ is free with exponents $(a,b,c)$ since it is boolean. Assuming that $(\mathscr{D},\mu)$ is free, the arrangement has exponents equal to one of $(a+3,b,c)_{\leq},(a+2,b+1,c)_{\leq},\dots,(a,b,c+3)_{\leq}$, thanks to [HRS17, Rem. 2.8]. Writing $E:=ab+ac+bc$ this gives $\operatorname{GMP}(2)$ as follows:
Exponents $(a+3,b,c)_{\leq}$ $(a,b+3,c)_{\leq}$ $(a,b,c+3)_{\leq}$ $(a+2,b+1,c)_{\leq}$ $(a+2,b,c+1)_{\leq}$
@@
By [CK24, Prop. 3.1], ${\mathscr{A}}_{C_{3}}$ is free with exponents $(1,3,3)$. Let $H=H_{123}$ and note that ${\mathscr{A}}_{C_{3}}^{H}$ is free with exponents $(1,2)$, since $|{\mathscr{A}}_{C_{3}}^{H}|=3$. According to Theorem 2.10, ${\mathscr{A}}_{C_{3}}\backslash\{H\}=\mathscr{D}$ is not free. By Theorem 2.25, $(\mathscr{D},\mu)$ is free if and only if ${\mathscr{E}}(\mathscr{D},\mu)$ free. However, ${\mathscr{E}}(\mathscr{D},\mu)_{X}=\mathscr{D}$ for $X=H_{1}\cap H_{2}\cap H_{3}$. So ${\mathscr{E}}(\mathscr{D},\mu)$ cannot be free, by Theorem 2.1. ∎
@@
Analogous to multiplicities, we occasionally denote by $\mu=(\mu(H_{1}),\mu(H_{2}),\dots,\mu(H_{n}))$, we write $\sigma$ as $\sigma=(\sigma(H_{1}),\sigma(H_{2}),\dots,\sigma(H_{n}))$. We use this notation in the following lemma.
@@
The idea of the proof is to increase the multiplicity of $H_{12}$ by $1$ which results in a multiplicity $\mu_{1}=(m_{1},m_{2},m_{3}+1)$ such that (2.26) is satisfied and thus ${\mathscr{E}}({\mathscr{A}},\mu_{1})$ is a free extension of $({\mathscr{A}},\mu_{1})$. Now we remove a hyperplane $H$ so that ${\mathscr{E}}({\mathscr{A}},\mu_{1})\backslash\{H\}={\mathscr{E}}_{\sigma}({\mathscr{A}},\mu)$ is free, thanks to Theorem 2.10. We demonstrate this in one case while the other cases are handled analogously.
For (i) let $\mu_{1}=(m_{1},m_{2},m_{3}+1)=(even,even,even)$. The extended arrangement ${\mathscr{A}}_{1}:={\mathscr{E}}_{0}({\mathscr{A}},\mu_{1})$ is free, by Theorem 2.25. Note that for $\sigma=(0,0,1),H=\ker(x_{1}+x_{2}+(\frac{m_{3}-1}{2})x_{3})$ and ${\mathscr{A}}_{2}:={\mathscr{E}}_{\sigma}({\mathscr{A}},\mu)$, we have ${\mathscr{A}}_{2}={\mathscr{A}}_{1}\backslash\{H\}$. So we start at the free arrangement ${\mathscr{A}}_{1}$, fix the hyperplane $H$, and use Theorem 2.10 to derive the freeness of ${\mathscr{A}}_{2}$. Note that the exponents of ${\mathscr{A}}_{1}$ can be derived by using Theorems 2.16 and 2.2. The following case by case investigation is necessary due to [Y09, Lem. 3.1].
@@
even even even $1$ $\geq 2$ $\geq 2$ $\sigma(H_{i})=0$ and $\sigma(H_{ij})=1\iff\mu(H_{ij})$ odd $(b)$
even even even $1$ $1$ $\geq 2$ $\sigma(H_{i})=0$ and $\sigma(H_{ij})=1\iff\mu(H_{ij})$ odd $(a)$
@@
Deleting a hyperplane from ${\mathscr{A}}_{C_{3}}$ results in either ${\mathscr{A}}_{P_{3}}={\mathscr{A}}(A_{3})$ or in $\mathscr{D}$. All free multiplicities on the reflection arrangement ${\mathscr{A}}(A_{3})$ have been classified, see [ATW08], [ANN09], [DFMS20]. Since all free multiplicities are known on both types of deletions on ${\mathscr{A}}_{C_{3}}$, the next step is to classify all free multiplicities on ${\mathscr{A}}_{C_{3}}$ itself.
@@ 5.3. Some free multiplicities on ${\mathscr{A}}_{C_{3}}$
Let $H=H_{123}$. Owing to Theorem 5.9, we see that $({\mathscr{A}}_{C_{3}}^{\prime},\mu^{\prime})$ is free with exponents $(2k+1,2k+1,2k+1)$. Now restrict $({\mathscr{A}}_{C_{3}},\mu)$ to $H$ and calculate $\mu^{*}$ by using Theorem 2.16 on the following rank $2$ flats:
@@
Theorem 5.12 ([ATW08, Thm. 5.10]).
Suppose ${\mathscr{A}}$ is supersolvable with a filtration ${\mathscr{A}}={\mathscr{A}}_{r}\supset{\mathscr{A}}_{r-1}\supset\dots\supset{\mathscr{A}}_{2}\supset{\mathscr{A}}_{1}$ and $r\geq 2$ as in Definition 5.11. Let $\mu$ be a multiplicity on ${\mathscr{A}}$ and let $\mu_{i}=\mu|_{{\mathscr{A}}_{i}}$. Let $\exp({\mathscr{A}}_{2},\mu_{2})=(d_{1},d_{2},0,\dots,0)$. Suppose that for each $H^{\prime}\in{\mathscr{A}}_{d}\backslash{\mathscr{A}}_{d-1}$, $H^{\prime\prime}\in{\mathscr{A}}_{d-1}$ $(d=3,\dots,r)$ and $X:=H^{\prime}\cap H^{\prime\prime}$, we either have
@@
Then $({\mathscr{A}},\mu)$ is inductively free with $\exp({\mathscr{A}},\mu)=(d_{1},d_{2},|\mu_{3}|-|\mu_{2}|,\dots,|\mu_{r}|-|\mu_{r-1}|,0,\dots,0)$.
@@
with $\mu^{*}=(k,k,r,1)$. If $r\geq 2k$, then the restriction is free with exponents $(2k+1,r)$ since deleting the hyperplane with multiplicity $1$ gives a free arrangement with exponents $(2k,r)$ and using Theorem 2.10 shows that $({\mathscr{A}}_{C_{3}},\mu)$ is inductively free with exponents $(r,2k+1,2k+1)$. Deleting $H_{12}$ with Theorem 2.10 and Lemma 5.3 until $r=2k-5$ (and we end up at exponents $(2k-1,2k-1,2k-1)$) shows the other claim. In particular: If $r=3$, then $\mu=(3,3,3,1,1,1,3)$ and $({\mathscr{A}}_{C_{3}},\mu)$ is free with exponents $\exp({\mathscr{A}}_{C_{3}},\mu)=(5,5,5)$. Restricting to an arbitrary $H\in{\mathscr{A}}_{C_{3}}$ results in $|({\mathscr{A}}_{C_{3}},\mu^{*})|=8$. So Remark 2.11 shows that $({\mathscr{A}}_{C_{3}}^{\prime},\mu^{\prime})$ can never be free and therefore $({\mathscr{A}}_{C_{3}},\mu)$ is neither inductively nor additively free. ∎
@@
Acknowledgments: The authors acknowledge the financial support of the DFG-priority program SPP 2458 “Combinatorial Synergies” (Grant #RO 1072/25-1 (project number: 539865068) to G. Röhrle). We are grateful to M. DiPasquale, T. Hoge, and M. Wakefield for some helpful comments.
@@ Appendix A Induction and Recursion tables
Table 2. Inductive chain for $({\mathscr{A}}_{G_{1}},\mu_{1})$ for $\mu=(2,2,2,1,2,1,1,1,2,1,1,1,1,1)$
@@
Table 3. Inductive chain for $({\mathscr{A}}_{C_{3}},\mu)$ for $\mu\equiv 3$
@@
Table 4. Recursive chain for $({\mathscr{A}}_{G_{2}},\mu_{2})$

### v3

@@ 5. Classifying free multiplicities on connected subgraph arrangements
Since connected subgraph arrangements are locally $A_{2}$ and admit a positive system of defining equations, the extension techniques introduced by Yoshinaga in [18] are well suited to attack the problem of classifying free multiplicities on connected subgraph arrangements and their subarrangements.
We define a subarrangement of ${\mathscr{A}}_{C_{3}}$ for which we give a complete list of free multiplicities. This classification was achieved previously by DiPasquale and Wakefield in [7], where this arrangement appeared in a different context, using homological algebra methods.
@@
[…]
@@
    If $\mu=(k,1,k,2k-4)$ or $\mu=(k,2k-4,k,1)$, then $({\mathscr{A}},\mu)$ is free with exponents (2k2,2k1)(2k-2,2k-1).
@@
    If $\mu=(k+1,1,k+1,2k-4)$ or $\mu=(k+1,2k-4,k+1,1)$, then $({\mathscr{A}},\mu)$ is free with exponents (2k1,2k)(2k-1,2k).
@@
(3) and (4): Start with $\mu^{\prime}=(k,0,k,2k-4)$ resp. $\mu^{\prime}=(k+1,0,k+1,2k-4)$ and use Theorem 2.16 to see that $({\mathscr{A}},\mu^{\prime})$ is free with exponents (2k2,2k2)(2k-2,2k-2) resp. (2k1,2k1)(2k-1,2k-1). So $({\mathscr{A}},\mu)$ has to be free with exponents (2k2,2k1)(2k-2,2k-1) resp. (2k1,2k)(2k-1,2k).
@@ 5.2. Non-Free multiplicities on $\mathscr{D}$
For the forward implication assume that $(\mathscr{D},\mu)$ is free. For some $X\in L({\mathscr{A}})_{2}$ and the corresponding rank $2$ localization ${\mathscr{A}}_{X}=\{H_{1},H_{2},H_{3}\}$ or ${\mathscr{A}}_{X}=\{H_{1},H_{2}\}$ which is free with $\exp({\mathscr{A}}_{X},\mu_{X})=(d_{1}^{X},d_{2}^{X})$, we write its summand $d_{1}^{X}d_{2}^{X}$ in $\operatorname{LMP}(2)$ as $\mu(H_{1}),\mu(H_{2}),\mu(H_{3})]$ or $\mu(H_{1}),\mu(H_{2})]$ respectively. Using this notation, $\operatorname{LMP}(2)$ of $(\mathscr{D},\mu)$ equals
@@
a(b+1)+a(c+1)+b(c+1)+a+b+c+3=3a+2b+c+ab+ac+bc+3.a(b+1)+a(c+1)+b(c+1)+a+b+c+3=3a+2b+c+ab+ac+bc+3.
Define the multiplicity $\mu_{1}=(a,b,c,0,0,0)$ on $\mathscr{D}$, then $(\mathscr{D},\mu_{1})$ is free with exponents $(a,b,c)$ since it is boolean. Assuming that $(\mathscr{D},\mu)$ is free, the arrangement has exponents equal to one of $(a+3,b,c)_{\leq},(a+2,b+1,c)_{\leq},\dots,(a,b,c+3)_{\leq}$, thanks to [10, Rem. 2.8]. Writing $E:=ab+ac+bc$ this gives $\operatorname{GMP}(2)$ as follows:
Exponents $_{\leq} (a,b+3,c)(a,b+3,c)_{\leq}$ $(a,b,c+3)_{\leq}$ $_{\leq} (a+2,b,c+1)(a+2,b,c+1)_{\leq}$
@@
[…]
@@
We occasionally denote a multiplicity $\mu$ by $\mu=(\mu(H_{1}),\mu(H_{2}),\dots,\mu(H_{n}))$, likewise for $\sigma$. We use this notation in the following lemma.
@@
[…]
@@
even even even 11 $\geq 2$ $\geq 2$ $\sigma(H_{i})=0$ and $\sigma(H_{ij})=1\iff\mu(H_{ij})$ odd $(b)$
even even even 11 11 $\geq 2$ $\sigma(H_{i})=0$ and $\sigma(H_{ij})=1\iff\mu(H_{ij})$ odd $(a)$
@@
Deleting a hyperplane from ${\mathscr{A}}_{C_{3}}$ results in either ${\mathscr{A}}_{P_{3}}={\mathscr{A}}(A_{3})$ or in $\mathscr{D}$. All free multiplicities on the reflection arrangement ${\mathscr{A}}(A_{3})$ have been classified, see [4], [2], [6]. Since all free multiplicities are known on both types of deletions on ${\mathscr{A}}_{C_{3}}$, the next step is to classify all free multiplicities on ${\mathscr{A}}_{C_{3}}$ itself.
@@ 5.3. Some free multiplicities on ${\mathscr{A}}_{C_{3}}$
Let $H=H_{123}$. Owing to Theorem 5.9, we see that $({\mathscr{A}}_{C_{3}}^{\prime},\mu^{\prime})$ is free with exponents $(2k+1,2k+1,2k+1)$. Now restrict $({\mathscr{A}}_{C_{3}},\mu)$ to $H$ and calculate $\mu^{*}$ by using Theorem 2.16 on the following rank 22 flats:
@@
Theorem 5.12 ([4, Thm. 5.10]).
Suppose $\mathscr{A}$} is supersolvable with a filtration ${\mathscr{A}}={\mathscr{A}}_{r}\supset{\mathscr{A}}_{r-1}\supset\dots\supset{\mathscr{A}}_{2}\supset{\mathscr{A}}_{1}$ and $r\geq 2$ as in Definition 5.11. Let $\mu$ be a multiplicity on $\mathscr{A}$} and let $\mu_{i}=\mu|_{{\mathscr{A}}_{i}}$. Let $\exp({\mathscr{A}}_{2},\mu_{2})=(e_{1},e_{2},0,\dots,0)$. Suppose that for each $H^{\prime}\in{\mathscr{A}}_{d}\backslash{\mathscr{A}}_{d-1}$, $H^{\prime\prime}\in{\mathscr{A}}_{d-1}$ $(d=3,\dots,r)$ and $X:=H^{\prime}\cap H^{\prime\prime}$, we either have
@@
Then $({\mathscr{A}},\mu)$ is inductively free with $\exp({\mathscr{A}},\mu)=(e_{1},e_{2},|\mu_{3}|-|\mu_{2}|,\dots,|\mu_{r}|-|\mu_{r-1}|,0,\dots,0)$.
@@
with $\mu^{*}=(k,k,r,1)$. If $r\geq 2k$, then the restriction is free with exponents $(2k+1,r)$ since deleting the hyperplane with multiplicity $1$ gives a free arrangement with exponents $(2k,r)$ and using Theorem 2.10 shows that $({\mathscr{A}}_{C_{3}},\mu)$ is inductively free with exponents $(r,2k+1,2k+1)$. Deleting $H_{12}$ with Theorem 2.10 and Lemma 5.3 until r=2k5r=2k-5 (and we end up at exponents (2k1,2k1,2k1)(2k-1,2k-1,2k-1)) shows the other claim. In particular: If $r=3$, then $\mu=(3,3,3,1,1,1,3)$ and $({\mathscr{A}}_{C_{3}},\mu)$ is free with exponents $\exp({\mathscr{A}}_{C_{3}},\mu)=(5,5,5)$. Restricting to an arbitrary $H\in{\mathscr{A}}_{C_{3}}$ results in $|({\mathscr{A}}_{C_{3}},\mu^{*})|=8$. So Remark 2.11 shows that $({\mathscr{A}}_{C_{3}}^{\prime},\mu^{\prime})$ can never be free and therefore $({\mathscr{A}}_{C_{3}},\mu)$ is neither inductively nor additively free. ∎
@@
Acknowledgments: The authors acknowledge the financial support of the DFG-priority program SPP 2458 “Combinatorial Synergies” (Grant #RO 1072/25-1 (project number: 539865068) to G. Röhrle). We are grateful to M. DiPasquale, T. Hoge, and M. Wakefield for some helpful comments. We would also like to thank the anonymous referee for pointing out some improvements in the exposition.
@@ Appendix A Induction and Recursion tables
Table 2. Inductive chain for $({\mathscr{A}}_{G_{1}},\mu_{1})$ for $\mu=(2,2,2,1,2,1,1,1,2,1,1,1,1,1)$.
@@
Table 3. Inductive chain for $({\mathscr{A}}_{C_{3}},\mu)$ for $\mu\equiv 3$.
@@
Table 4. Recursive chain for $({\mathscr{A}}_{G_{2}},\mu_{2})$.